OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n. The integer n^3 + 2n is divisible by 3 for every positive integer n.

OpenStudy (primeralph):

By induction. Just say since it happens for n consecutive numbers, it's happens for all other numbers.

OpenStudy (anonymous):

1. for n=1 statement is true, we can check it: 1+2=3 => 3 divides 3 2. Assume that the statement is true for n=k: 3 divides k^3+2k 3. Then it has to be right for n=k+1 - let's check (k+1)^3+2(k+1)=k^3+3k^2+5k+3=(k^3+2k) + (3k^2+3k+3) = (k^3+2k) + 3(k^2+k+1) - we can divide this by 3, so the statement is proved

OpenStudy (anonymous):

done thank yoU!!

OpenStudy (primeralph):

What happened to my medal?

OpenStudy (anonymous):

No problem :)

OpenStudy (blurbendy):

mathematical induction, you so sexy

OpenStudy (primeralph):

@blurbendy Ahem!

OpenStudy (blurbendy):

hehe

OpenStudy (anonymous):

Let P(n)=n^3+2n for n=1,1^{3}+2*1=3 which is divisible by 3 Hence P(1) is true. Assume that P(k) is true. P(k)=k^3+2k is divisible by 3 k^3+2k=3l where l is an integer. k^3=3l-2k P(k+1)=(k+1)^3+2(k+1) \[=k ^{3}+1^{3}+3k*1\left( k+1 \right)+2k+2\] \[=k ^{3}+1+3k ^{2}+3k+2k+2\] \[=k ^{3}+2k+3k ^{2}+3k+3=3l-2k+2k+3\left( k ^{2} +k+1\right)\] \[=3\left( l+k ^{2}+k+1 \right)=3*an integer.\] Hence P(k+1) is true . Hence by induction P(n)is true forall n

OpenStudy (anonymous):

i like that response better lol